A 0.5 m length of current-carrying wire kept perpendicular to a magnetic field of
magnitude 200 mT experiences a force of 1.5 mN. What is the current flowing in the
wire.
1
Expert's answer
2020-03-03T10:06:55-0500


Ampere force acts on a conductor with current in a magnetic field

"F=B \cdot I \cdot l"

where will we write

"I=\frac {F}{B \cdot l}=\frac {1.5}{200 \cdot 0.5}=15 mA"